OpenStudy (anonymous):

A family went to Hershey Park and paid $12 for each adult and $8 for each child. A group of 15 people went to the park and it cost $160. This system of equations models this situation where x is the number of adults and y is the number of children. x + y = 15 12x + 8y = 160 How many adults were in the group?

OpenStudy (anonymous):

pleaseee helppp!!! i will medal.

OpenStudy (tkhunny):

Have you considered solving the system?

OpenStudy (anonymous):

i tried i just dont understand.

OpenStudy (tkhunny):

Simplify. Divide the second equation by 4. What do you get?

OpenStudy (anonymous):

if i divide 12 by 4 i get 3 and if i divide 8 by 4 its 2. But once i put them in i got 12(3)+8(2)=160 which is not true

OpenStudy (anonymous):

ok audrey it is a simple equation :)

OpenStudy (anonymous):

12x + 8(15-x) = 120

OpenStudy (tkhunny):

Why did you do that? 12x + 8y = 160 Divide by 4 3x + 2y = 40 That's all. Don't try to solve up front. Just take one step at a time. Make sense?

OpenStudy (anonymous):

can you see where my equation comes from or do you need me to explain?

OpenStudy (anonymous):

i got it. i will tell you my answer once i solve.

OpenStudy (anonymous):

oh shoot 160 not 120 my mistake

OpenStudy (anonymous):

i copied it wrong just replace the 120 and 160 solve for x and that will be the number of adults

OpenStudy (anonymous):

i feel so stupid, how do i solve the equation?

OpenStudy (anonymous):

ok dont worry :) , so we have 12x + 8(15-x) = 160 12x -8x + 120 = 160 4x = 40 x=10

OpenStudy (anonymous):

does that make sense? and if we do 10(12) we get 120 + 8(5) for the kids equals 160

OpenStudy (anonymous):

yes, thank you so much!!

OpenStudy (anonymous):

happy to help audrey! It makes sense right?

OpenStudy (tkhunny):

Too bad we don't know if YOU can do it. @MathMan77 did all the work. Why don't you show us one.

OpenStudy (anonymous):

good idea!

OpenStudy (anonymous):

yes, it makes sence now. what equation?

OpenStudy (anonymous):

I will make one up adults cost 10 kids cost 5 we have a group of 6 that spend 50 dollars can you give me the equation?

OpenStudy (anonymous):

x + y = 6 10x + 5y = 50 These were supplied in problem above

OpenStudy (tkhunny):

Please simplify first. It will make your life easier. Divide the second equation by 5. Don't substitute ANYTHING until you have done that.

OpenStudy (anonymous):

25? because 10/5 is 2 and 5/5 would cancel, so 2=50 then divide that and get 25??

OpenStudy (anonymous):

heres what i do: we know x is number of adults y is kids and total is 6 so, x = adults y(kids) = total(6) - adults(x) so y = 6 - x we put that into the bottom eq with substitution and get...

OpenStudy (anonymous):

how do we sub that in? :(

OpenStudy (anonymous):

substitute y with (6-x) and distribute